# Rethinking Approximate Gaussian Inference in Classification

### v2

@@ E.1 Predictive Formulas
We gather formulas for all predictive estimators $\bm{\hat{p}}$ of the true predictive $\mathbb{E}_{\bm{P}\sim\bm{a}_{*}\mathcal{N}(\bm{\mu},\bm{\Sigma})}[\bm{P}]$, $\bm{\Sigma}=\bm{\operatorname{\mathbf{diag}}}(\bm{\sigma}^{2})$, used in our experiments (Section 6).
@@ E.1.1 Softmax
Mean Field (Section A.4)
The Mean Field predictive [21] uses the following approximation for the true predictive:
@@
The Laplace Bridge predictive [14] approximates the true predictive as follows:
@@ E.1.2 Exp
Our closed-form predictive for the exp activation function (Section 2.1) is given by
@@ E.1.3 NormCDF
The closed-form predictive for the normCDF activation function (Section 2.2) is computed as
@@ E.1.4 Sigmoid
For the sigmoid activation function (Section 2.2), the closed-form predictive can be computed as
@@ E.2.1 Softmax
The Laplace Bridge method [14] uses the following Dirichlet parameters:
@@ E.2.2 Exp
Exp uses the closed-form parameters derived from Moment Matching the Gaussian pushforwards (Section 3.1):
@@ E.2.3 NormCDF
NormCDF uses the closed-form parameters derived from Moment Matching the Gaussian pushforwards (Section 3.1):
@@ E.2.4 Sigmoid
Sigmoid uses the closed-form parameters derived from Moment Matching the Gaussian pushforwards (Section 3.1):
@@ Appendix F List of Uncertainty Estimators
In this section, we list the uncertainty estimators used in our experiments (Section 6).
@@ F.2 Monte Carlo
Given $S$ Monte Carlo samples $\bm{\hat{p}}^{(1)},\dots,\bm{\hat{p}}^{(S)}$ with mean $\bm{\hat{p}}$, one can calculate a predictive as their average and derive the estimators in Section F.1. However, Monte Carlo samples allow one to calculate two additional estimators detailed below.
@@ F.3 Dirichlet
$-\sum_{c=1}^{C}\frac{\gamma_{c}}{\sum_{c^{\prime}=1}^{C}\gamma_{c^{\prime}}}\left(\psi(\gamma_{c}+1)-\psi\left(\sum_{c^{\prime}=1}^{C}\gamma_{c^{\prime}}+1\right)\right)$ (83)
@@ Appendix G Experimental Setup
To answer the first question, we evaluate our closed-form predictives (Sigmoid, NormCDF, Exp) and moment-matched Dirichlet distributions against softmax models equipped with approximate inference tools (Laplace Bridge [14], Mean Field [21], Monte Carlo sampling). We consider Heteroscedastic Classifiers (HET) [5], Spectral-Normalized Gaussian Processes (SNGP) [20], and last-layer Laplace approximation methods [8] as backbones (see Appendix I for details).
The resulting 21 (method, activation, predictive) triplets are evaluated on ImageNet-1k [9] and CIFAR-10 [17] on five metrics aligning with practical needs from uncertainty estimates [26]:
@@
See Appendix H for details.
For ImageNet, we treat ImageNet-C [13] samples with 15 corruption types and 5 severity levels as OOD samples. For CIFAR-10, we use the CIFAR-10C corruptions.
@@
For ImageNet evaluation, we use a ResNet-50 backbone pretrained with the softmax activation function, and train each (method, activation) pair for 50 ImageNet-1k epochs following Mucsányi et al. [27]. We train with the LAMB optimiser [40] using a batch size of $128$ and gradient accumulation across $16$ batches, resulting in an effective batch size of $2048$, following Tran et al. [36]. We further use a cosine learning rate schedule with a single warmup epoch using a warmup learning rate of $0.0001$. The learning rate is treated as a hyperparameter and selected from the interval $[0.0005,0.05]$ based on the validation performance. The weight decay is selected from the set $\{0.01,0.02\}$. During training, we keep track of the best-performing checkpoint on the validation set and load it before testing. We search for ideal hyperparameters with a ten-step Bayesian Optimization scheme [33] in Weights & Biases [3] based on the negative log-likelihood.
On CIFAR-10, we train ResNet-28 models from scratch for 100 epochs. The only exceptions are the SNGP models that are trained for 125 epochs [20]. We train with Momentum SGD using a batch size of $128$ and no gradient accumulation. Similarly to ImageNet, we use a cosine learning rate schedule but with five warmup epochs and warmup learning rate $1\mathrm{e}{-5}$. The learning rate is also treated as a hyperparameter on CIFAR-10. We use the interval $[0.05,1]$ for Sigmoid and NormCDF, and $[0.01,0.15]$ for Softmax and Exp. The optimal learning rates for Sigmoid and NormCDF are generally larger, as the class-wise binary cross-entropies are averaged instead of summed. The weight decay is selected from the interval $[1\mathrm{e}{-6},1\mathrm{e}{-4}]$. Similarly to ImageNet, we use the best-performing checkpoint in the tests and use a ten-step Bayesian Optimization scheme to select performant hyperparameters.
@@
Appendix H Benchmark Metrics
@@
H.1 Log Probability Proper Scoring Rule for the Predictive
First, we briefly discuss proper and strictly proper scoring rules over general probability measures based on [26].
@@ H.1 Log Probability Proper Scoring Rule for the Predictive
H.2 Expected Calibration Error
To set up the required quantities for the Expected Calibration Error (ECE) metric [29], we follow the steps below, based on [26].
@@ H.2 Expected Calibration Error
The Expected Calibration Error (ECE) metric [29] is then defined as
@@
H.3 Binary Log Probability Proper Scoring Rule for Correctness Prediction
@@
$S(\tilde{c},\ell)=\begin{cases}\log c&\mathrm{if}\ \ell=1\\ \log(1-\tilde{c})&\mathrm{if}\ \ell=0\end{cases}=\ell\log\tilde{c}+(1-\ell)\log(1-\tilde{c}).$ (93)
One can show that this is indeed a strictly proper scoring rule [26].
H.4 Accuracy
@@
H.5 Area Under the Receiver Operating Characteristic Curve for Out-of-Distribution Detection
@@
Given uncertainty estimates $u_{n}\equiv u(x_{n})$ and target binary labels $o_{i}$ on a balanced dataset $(x_{n},o_{n})_{n=1}^{N}$, as well as a threshold $t\in\mathbb{R}$, we predict $1$ (out-of-distribution) when $u_{n}\geq t$ and 00 (in-distribution) when $u_{n}<t$. This lets us define the following index sets:
[…]
@@
One can draw a curve of $(\mathrm{FPR}(t),\mathrm{TPR}(t))$ for all $t$ from $\infty$- ∞ to +$\infty$+ ∞. This is the ROC curve. The area under this curve quantifies how well the uncertainty estimator $u(x)$ can separate in-distribution and out-of-distribution inputs.
Appendix I Benchmarked Methods
@@
I.1 Spectral Normalized Gaussian Process
Spectral normalized Gaussian processes (SNGP) [20] use spectral normalization of the parameter tensors for distance-awareness and a last-layer Gaussian process approximated by Fourier features to capture uncertainty. For an input $x\in\mathcal{X}$ and number of classes $C$, they predict a $C$-variate Gaussian distribution
@@ I.1 Spectral Normalized Gaussian Process
The method applies spectral normalization to the hidden weights in each layer using a power iteration scheme with a single iteration per batch to obtain the largest singular value. Liu et al. [20] claim this helps with input distance awareness.
I.2 Heteroscedastic Classifier
Heteroscedastic classifiers (HET) [5] construct a Gaussian distribution in the logit space to model per-input uncertainties:
@@ I.2 Heteroscedastic Classifier
To learn the per-input covariance matrices from the training set, one has to construct a predictive estimate from $\mathcal{N}(\bm{f}(x),\bm{\Sigma}(x))$ using any of the methods in Appendix E. This predictive estimate is then trained using a standard cross-entropy (NLL) loss.
@@
The off-diagonal terms of the covariance matrix do not affect the approximate predictive (Eq. 12). This means that, in our framework, one can discard the low-rank term $\bm{V}(x)$ and only model the diagonal term $\bm{d}(x)$ without a decrease in expressivity. To keep comparisons fair and use the same backbone with the same number of parameters, we also only model $\bm{d}(x)$ for softmax-based predictives.
I.3 Laplace Approximation
The Laplace approximation [8] approximates the posterior $p(\bm{\theta}\mid\mathcal{D})$ over the network parameters $\bm{\theta}$ for a Gaussian prior $p(\bm{\theta})$ and likelihood defined by the network architecture by a Gaussian. In its simplest form, it uses the maximum a posteriori (MAP) weights $\bm{\theta}_{\text{MAP}}\in\mathbb{R}^{P}$ as the mean and the inverse Hessian of the regularised loss over the training set ~$\tilde{\mathcal{L}}(\bm{\theta};\mathcal{D})=\mathcal{L}(\bm{\theta};\mathcal{D})+\lambda\|\bm{\theta}\|_{2}^{2}$ evaluated at the MAP as the covariance matrix:
@@ I.3 Laplace Approximation
This is a locally optimal post-hoc Gaussian approximation of the true posterior $p(\bm{\theta}\mid\mathcal{D})$ based on a second-order Taylor approximation. For details, see [35].
For deep neural networks, the Hessian matrix is often replaced with the Generalized Gauss-Newton (GGN) matrix. The GGN is guaranteed to be positive semidefinite even for suboptimal weights and has efficient approximation schemes, such as Kronecker-Factored Approximate Curvature [24] or low-rank approximations.
@@
We use a last-layer KFAC Laplace variant in our experiments and use the full training set for calculating the GGN instead of a mini-batch based on recent works on the bias in mini-batch estimates [35].
Appendix J CIFAR-10 Experiments
This appendix section repeats the experiments presented in the main paper on the CIFAR-10 dataset. For a detailed description of the experimental setup, refer to Appendix G. Appendix H describes the used tasks and metrics.
As stated in the main paper, our two research questions are:
  • What are the effects of changing the learning objective? (Section J.2)
  • Do we have to sacrifice performance for sample-free predictives? (Section J.1)
J.1 Quality of Sample-Free Predictives
Table J.1: Comparison of ECE results for different predictives using a fixed Laplace backbone.
Method Mean $\pm$ 2 std
Softmax Laplace
MC 100 0.0096 $\pm$ 0.0013
MC 1000 0.0102 $\pm$ 0.0016
MC 10 0.0120 $\pm$ 0.0013
Mean Field 0.0121 $\pm$ 0.0029
Laplace Bridge Predictive 0.5933 $\pm$ 0.0105
NormCDF Laplace
Closed-form 0.0074 $\pm$ 0.0012
MC 1000 0.0092 $\pm$ 0.0022
MC 100 0.0095 $\pm$ 0.0015
MC 10 $0$$\pm$ 0.0020
Similarly to the main paper, in this section, we investigate our first research question: whether there is a price to pay for sample-free predictives. Table J.1 showcases the two best-performing (activation, method) pairs on the ECE metric and the CIFAR-10 dataset: Softmax and NormCDF Laplace. Mean Field (MF) is a strong alternative for sample-free predictives, but it has no guarantees and can fall behind MC sampling (see also Fig. 1). Empirically, our closed-form predictives always perform on par with MC sampling.
J.2 Effects of Changing the Learning Objective
As in the main paper, in this section, we use the best-performing predictive and estimator (see Appendix E) for softmax models and employ our methods with the closed-form predictives.
J.2.1 Calibration and Proper Scoring
We first evaluate calibration using the log probability scoring rule [12] and the Expected Calibration Error (ECE) metric [29]. Fig. 3(b) shows that on CIFAR-10, the score of our closed-form predictives (Sigmoid, NormCDF) are consistently better than the corresponding softmax results for all methods.
Figure J.1: CIFAR-10 ECE results. Our closed-form predictives ( ) outperform Softmax ( ) on HET and SNGP. Laplace tunes its hyperparameters based on the ECE metric – NormCDF Laplace is the overall best method. Note the restricted $y$-limits for readability.
Fig. J.1 shows that on CIFAR-10, our closed-form predictives have a clear advantage on HET and SNGP. Laplace is a post-hoc method that tunes its hyperparameters on the ECE metric, hence its enhanced performance. Our NormCDF predictive is on par with Softmax.
J.2.2 Correctness Prediction
Figure J.2: CIFAR-10 log probability proper scoring results for the binary correctness prediction task. Our closed-form predictives ( ) consistently outperform Softmax ( ) on all methods.
Fig. J.2 shows that on the correctness prediction task, our closed-form predictives outperform all Softmax predictives across all methods, as measured by the log probability proper scoring rule.
J.2.3 Accuracy
Figure J.3: Closed-form predictives do not sacrifice accuracy. CIFAR-10 accuracies. Our closed-form predictives ( ) either outperform or are on par with Softmax ( ) across all methods.
Closed-form predictives do not sacrifice accuracy. Fig. J.3 evidences this claim on CIFAR-10: our closed-form predictives either outperform or are on par with Softmax predictives. The most accurate method is Sigmoid HET. These results support the findings of Wightman et al. [38] that showcase desirable training dynamics of the class-wise cross-entropy loss.
J.2.4 Out-of-Distribution Detection
Figure J.4: CIFAR-10C OOD detection AUROC results for severity level one. Across all methods, the best-performing predictive is closed-form ( ).
Finally, we consider the OOD detection task on a balanced mixture of ID (CIFAR-10) and OOD inputs. As OOD inputs, we consider corrupted CIFAR-10C samples. We use the AUROC metric to evaluate the methods’ performance. As shown in Fig. J.4, the best-performing method is Sigmoid SNGP, a closed-form method. Generally, Softmax performs on par with our closed-form predictives. Intuitively, separating ID and OOD samples does not require a fine-grained representation of uncertainty, unlike the ECE or proper scoring rules. Nevertheless, the closed-form predictives and second-order Dirichlet distributions are considerably cheaper to calculate than Softmax MC predictions (see Section 5).
@@
One can directly apply Eq. 10 to a neural network trained with the softmax cross-entropy loss in Eq. 4 to obtain closed-form predictives. However, we empirically found this approach to decrease performance, as the denominator of Eq. 10 increased to the order of billions during training. Considering a multivariate normal random variable $\bm{x}\sim\mathcal{N}(\bm{\mu},\Sigma)$ representing a logit distribution,
@@
To mitigate this issue, one may optimize the regularised cross-entropy loss
@@
for an appropriately tuned $\lambda$ hyperparameter. The latter formulation can be stably trained even at an ImageNet scale. However, there are fundamental limitations to the predictive given by Eq. 10. Namely, for isotropic logit-space Gaussian distributions, $\nicefrac{{\sigma^{2}(x)}}{{2}}$ introduces a constant shift in the logits, under which the softmax activation function is invariant. Therefore, the predictive approximation collapses into the MAP prediction. The SNGP method predicts such isotropic logit-space Gaussians. Further, we empirically found that the Laplace method’s predictives for models trained with the regularised cross-entropy loss were also approximately isotropic; thus, this integral approximation yielded diminishing returns.
For completeness, we share the ImageNet and CIFAR-10 results of $\varphi=\exp$ in Tables K.1 and K.2.
@@
Table K.3 shows results on ViT Little [11] backbones from the timm [37] library on the ImageNet validation set. As in the main paper, the hyperparameters are optimized using a ten-step Bayesian hyperparameter optimization scheme of Weights and Biases.

### v3

@@ E.1 Predictive Formulas
We gather formulas for all predictive estimators $\bm{\hat{p}}$ of the true predictive $\mathbb{E}_{\bm{P}\sim\bm{a}_{*}\mathcal{N}(\bm{\mu},\bm{\Sigma})}[\bm{P}]$, $\bm{\Sigma}=\bm{\operatorname{\mathbf{diag}}}(\bm{\sigma}^{2})$, used in our experiments (Section˜7).
@@ E.1.1 Softmax
Mean Field (Section˜A.4)
The Mean Field predictive [lu_mean-field_2021] uses the following approximation for the true predictive:
@@
The Laplace Bridge predictive [pmlr-v180-hobbhahn22a] approximates the true predictive as follows:
@@
Shekhovtsov & Flach.
The shekhovtsov_feed-forward_2018 approximation of the Gaussian-softmax integral yields:
$\bm{\hat{p}}=\frac{\bm{n}}{\bm{n}+\big(A-\bm{b}\big)^{s/\bm{s}}}$ (69)
where
$A:=\sum_{c=1}^{C}\exp(\tilde{\mu}_{c}/s),\;\bm{n}:=\exp(\bm{\tilde{\mu}}/\bm{s}),\;\bm{b}:=\exp(\bm{\tilde{\mu}}/s),$ (70)
$\bm{\tilde{\mu}}:=\bm{\mu}-\max_{c}\mu_{c}$ for numerical stability,
$\bm{\tau}^{2}:=\bm{\sigma}^{2}+G,\;\tau_{\text{min}}^{2}:=\min_{c}\tau_{c}^{2},\;s:=\sqrt{\tfrac{2\,\tau_{\text{min}}^{2}}{L}},\;\bm{s}:=\sqrt{\tfrac{\bm{\tau}^{2}+\tau_{\text{min}}^{2}}{L}}$ (71)
and $L:=\pi^{2}/3$, $G:=L/2$.
@@ E.1.2 Exp
Our closed-form predictive for the exp activation function (Section˜3.1) is given by
@@ E.1.3 NormCDF
The closed-form predictive for the normCDF activation function (Section˜3.2) is computed as
@@ E.1.4 Sigmoid
For the sigmoid activation function (Section˜3.2), the closed-form predictive can be computed as
@@ E.2.1 Softmax
The Laplace Bridge method [pmlr-v180-hobbhahn22a] uses the following Dirichlet parameters:
@@ E.2.2 Exp
Exp uses the closed-form parameters derived from Moment Matching the Gaussian pushforwards (Section˜4.1):
@@ E.2.3 NormCDF
NormCDF uses the closed-form parameters derived from Moment Matching the Gaussian pushforwards (Section˜4.1):
@@ E.2.4 Sigmoid
Sigmoid uses the closed-form parameters derived from Moment Matching the Gaussian pushforwards (Section˜4.1):
@@ Appendix F List of Uncertainty Estimators
In this section, we list the uncertainty estimators used in our experiments (Section˜7).
@@ F.2 Monte Carlo
Given $S$ Monte Carlo samples $\bm{\hat{p}}^{(1)},\dots,\bm{\hat{p}}^{(S)}$ with mean $\bm{\hat{p}}$, one can calculate a predictive as their average and derive the estimators in Section˜F.1. However, Monte Carlo samples allow one to calculate two additional estimators detailed below.
@@ F.3 Dirichlet
$\sum_{c=1}^{C}\frac{\gamma_{c}}{\sum_{c^{\prime}=1}^{C}\gamma_{c^{\prime}}}\left(\psi(\gamma_{c}+1)-\psi\left(\sum_{c^{\prime}=1}^{C}\gamma_{c^{\prime}}+1\right)\right)$ (87)
@@ Appendix G Experimental Setup
To answer the first question, we evaluate our closed-form predictives (Sigmoid, NormCDF, Exp) and moment-matched Dirichlet distributions against softmax models equipped with approximate inference tools (Laplace Bridge [pmlr-v180-hobbhahn22a], Mean Field [lu_mean-field_2021], Monte Carlo sampling). We consider Heteroscedastic Classifiers (HET) [collier2021correlated], Spectral-Normalised Gaussian Processes (SNGP) [liu2020simple], and last-layer Laplace approximation methods [daxberger2021laplace] as backbones (see Appendix˜J for details).
The resulting 21 (method, activation, predictive) triplets are evaluated on ImageNet-1k [deng2009imagenet] and CIFAR-10 [krizhevsky2009learning] on five metrics aligning with practical needs from uncertainty estimates [mucsanyi2023trustworthy]:
@@
See Appendix˜I for details.
For ImageNet, we treat ImageNet-C [hendrycks2019benchmarking] samples with 15 corruption types and 5 severity levels as OOD samples. For CIFAR-10, we use the CIFAR-10C corruptions.
@@
For ImageNet evaluation, we use a ResNet-50 backbone pretrained with the softmax activation function, and train each (method, activation) pair for 50 ImageNet-1k epochs following mucsányi2024benchmarking. We train with the LAMB optimiser [you2019large] using a batch size of $128$ and gradient accumulation across $16$ batches, resulting in an effective batch size of $2048$, following tran2022plex. We further use a cosine learning rate schedule with a single warmup epoch using a warmup learning rate of $0.0001$. The learning rate is treated as a hyperparameter and selected from the interval $[0.0005,0.05]$ based on the validation performance. The weight decay is selected from the set $\{0.01,0.02\}$. During training, we keep track of the best-performing checkpoint on the validation set and load it before testing. We search for ideal hyperparameters with a ten-step Bayesian Optimization scheme [shahriari2015taking] in Weights & Biases [wandb] based on the negative log-likelihood.
On CIFAR-10, we train ResNet-28 models from scratch for 100 epochs. The only exceptions are the SNGP models that are trained for 125 epochs [liu2020simple]. We train with Momentum SGD using a batch size of $128$ and no gradient accumulation. Similarly to ImageNet, we use a cosine learning rate schedule but with five warmup epochs and warmup learning rate $1\mathrm{e}{-5}$. The learning rate is also treated as a hyperparameter on CIFAR-10. We use the interval $[0.05,1]$ for Sigmoid and NormCDF, and $[0.01,0.15]$ for Softmax and Exp. The optimal learning rates for Sigmoid and NormCDF are generally larger, as the class-wise binary cross-entropies are averaged instead of summed. The weight decay is selected from the interval $1\mathrm{e}{-6},1\mathrm{e}{-4}]$. Similarly to ImageNet, we use the best-performing checkpoint in the tests and use a ten-step Bayesian Optimization scheme to select performant hyperparameters.
@@
Appendix H Runtime Overhead of Monte Carlo Sampling
We measure the time to obtain the predictives from the logit-space Gaussians compared to the time of a forward pass on an NVIDIA RTX 2080 Ti GPU (with similar results on Tesla A100 GPUs). As shown in Table˜H.1, ResNet-50 with 1000 samples shows 6-8% overhead depending on the batch size, confirming our claim in the main text. In contrast, the runtime cost of computing our closed-form predictives is negligible compared to a forward pass, especially for larger batch sizes. These results demonstrate that even with a diagonal covariance structure, MC sampling yields a nontrivial overhead, especially for large numbers of classes and samples. In comparison, our closed-form approach has negligible computational cost.
Table H.1: Runtime overhead of Monte Carlo sampling on ImageNet as a percentage of total inference time across models and batch sizes.
Batch Size Mean Field 1 sample 10 samples 100 samples 1000 samples
ResNet-26 (C=10)
1 2.75% 9.82% 10.63% 10.70% 10.70%
16 2.04% 10.61% 11.51% 11.51% 11.92%
32 2.11% 10.98% 11.94% 11.96% 12.42%
64 2.11% 10.88% 11.90% 11.95% 12.34%
128 2.04% 8.12% 8.84% 8.96% 9.61%
256 1.17% 4.31% 4.73% 4.93% 7.14%
Wide-ResNet-26-5 (C=100)
1 1.29% 8.54% 9.36% 9.40% 9.28%
16 1.16% 5.94% 6.58% 6.84% 8.56%
32 0.86% 3.36% 3.76% 3.98% 6.95%
64 0.47% 1.83% 2.03% 2.12% 5.82%
128 0.23% 0.90% 1.04% 1.04% 4.88%
256 0.12% 0.46% 0.53% 0.84% 4.47%
ResNet-50 (C=1000)
1 1.15% 5.43% 5.88% 6.09% 6.30%
16 0.32% 1.25% 1.43% 1.70% 8.04%
32 0.16% 0.64% 0.72% 1.34% 7.51%
64 0.09% 0.36% 0.41% 1.13% 7.75%
128 0.05% 0.20% 0.22% 1.00% 7.82%
256 0.03% 0.10% 0.18% 0.95% 8.05%
Appendix I Benchmark Metrics
@@
I.1 Log Probability Proper Scoring Rule for the Predictive
First, we briefly discuss proper and strictly proper scoring rules over general probability measures based on [mucsanyi2023trustworthy].
@@ I.1 Log Probability Proper Scoring Rule for the Predictive
I.2 Expected Calibration Error
To set up the required quantities for the Expected Calibration Error (ECE) metric [naeini2015obtaining], we follow the steps below, based on [mucsanyi2023trustworthy].
@@ I.2 Expected Calibration Error
The Expected Calibration Error (ECE) metric [naeini2015obtaining] is then defined as
@@
I.3 Binary Log Probability Proper Scoring Rule for Correctness Prediction
@@
[…]
One can show that this is indeed a strictly proper scoring rule [mucsanyi2023trustworthy].
I.4 Accuracy
@@
I.5 Area Under the Receiver Operating Characteristic Curve for Out-of-Distribution Detection
@@
Given uncertainty estimates $u_{n}\equiv u(x_{n})$ and target binary labels $o_{i}$ on a balanced dataset $(x_{n},o_{n})_{n=1}^{N}$, as well as a threshold $t\in\mathbb{R}$, we predict $1$ (out-of-distribution) when $u_{n}\geq t$ and 0 (in-distribution) when $u_{n}<t$. This lets us define the following index sets:
$\text{True positives: }\mathrm{TP}(t)$ $=\left\{n:o_{n}=1\land u_{n}\geq t\right\}$ (99)
$\text{False positives: }\mathrm{FP}(t)$ $=\left\{n:o_{n}=0\land u_{n}\geq t\right\}$
$\text{False negatives: }\mathrm{FN}(t)$ $=\left\{n:o_{n}=1\land u_{n}<t\right\}$
$\text{True negatives: }\mathrm{TN}(t)$ $=\left\{n:o_{n}=0\land u_{n}<t\right\}.$
@@
One can draw a curve of $(\mathrm{FPR}(t),\mathrm{TPR}(t))$ for all $t$ from -$\infty$ to $\infty$. This is the ROC curve. The area under this curve quantifies how well the uncertainty estimator $u(x)$ can separate in-distribution and out-of-distribution inputs.
Appendix J Benchmarked Methods
@@
J.1 Spectral Normalised Gaussian Process
Spectral normalised Gaussian processes (SNGP) [liu2020simple] use spectral normalization of the parameter tensors for distance-awareness and a last-layer Gaussian process approximated by Fourier features to capture uncertainty. For an input $x\in\mathcal{X}$ and number of classes $C$, they predict a $C$-variate Gaussian distribution
@@
The method applies spectral normalization to the hidden weights in each layer using a power iteration scheme with a single iteration per batch to obtain the largest singular value. liu2020simple claim this helps with input distance awareness.
J.2 Heteroscedastic Classifier
Heteroscedastic classifiers (HET) [collier2021correlated] construct a Gaussian distribution in the logit space to model per-input uncertainties:
@@
To learn the per-input covariance matrices from the training set, one has to construct a predictive estimate from $\mathcal{N}(\bm{f}(x),\bm{\Sigma}(x))$ using any of the methods in Appendix˜E. This predictive estimate is then trained using a standard cross-entropy (NLL) loss.
@@
The off-diagonal terms of the covariance matrix do not affect the approximate predictive (Eq.˜12). This means that, in our framework, one can discard the low-rank term $\bm{V}(x)$ and only model the diagonal term $\bm{d}(x)$ without a decrease in expressivity. To keep comparisons fair and use the same backbone with the same number of parameters, we also only model $\bm{d}(x)$ for softmax-based predictives.
J.3 Laplace Approximation
The Laplace approximation [daxberger2021laplace] approximates the posterior $p(\bm{\theta}\mid\mathcal{D})$ over the network parameters $\bm{\theta}$ for a Gaussian prior $p(\bm{\theta})$ and likelihood defined by the network architecture by a Gaussian. In its simplest form, it uses the maximum a posteriori (MAP) weights $\bm{\theta}_{\text{MAP}}\in\mathbb{R}^{P}$ as the mean and the inverse Hessian of the regularised loss over the training set $\tilde{\mathcal{L}}(\bm{\theta};\mathcal{D})=\mathcal{L}(\bm{\theta};\mathcal{D})+\lambda\|\bm{\theta}\|_{2}^{2}$ evaluated at the MAP as the covariance matrix:
@@
This is a locally optimal post-hoc Gaussian approximation of the true posterior $p(\bm{\theta}\mid\mathcal{D})$ based on a second-order Taylor approximation. For details, see [tatzel2025debiasing].
For deep neural networks, the Hessian matrix is often replaced with the Generalised Gauss-Newton (GGN) matrix. The GGN is guaranteed to be positive semidefinite even for suboptimal weights and has efficient approximation schemes, such as Kronecker-Factored Approximate Curvature [martens2015optimizing] or low-rank approximations.
@@
We use a last-layer KFAC Laplace variant in our experiments and use the full training set for calculating the GGN instead of a mini-batch based on recent works on the bias in mini-batch estimates [tatzel2025debiasing].
@@
One can directly apply Eq.˜10 to a neural network trained with the softmax cross-entropy loss in Eq.˜4 to obtain closed-form predictives. However, we empirically found this approach to decrease performance, as the denominator of Eq.˜10 increased to the order of billions during training. Considering a multivariate normal random variable $\bm{x}\sim\mathcal{N}(\bm{\mu},\Sigma)$ representing a logit distribution,
@@
To mitigate this issue, one may optimise the regularised cross-entropy loss
@@
for an appropriately tuned $\lambda$ hyperparameter. The latter formulation can be stably trained even at an ImageNet scale. However, there are fundamental limitations to the predictive given by Eq.˜10. Namely, for isotropic logit-space Gaussian distributions, $\nicefrac{{\sigma^{2}(x)}}{{2}}$ introduces a constant shift in the logits, under which the softmax activation function is invariant. Therefore, the predictive approximation collapses into the MAP prediction. The SNGP method predicts such isotropic logit-space Gaussians. Further, we empirically found that the Laplace method’s predictives for models trained with the regularised cross-entropy loss were also approximately isotropic; thus, this integral approximation yielded diminishing returns.
For completeness, we share the ImageNet and CIFAR-10 results of $\varphi=\exp$ in Tables˜K.1 and K.2.
@@
Table˜K.3 shows results on ViT Little [dosovitskiy2020image] backbones from the timm [rw2019timm] library on the ImageNet validation set. As in the main paper, the hyperparameters are optimised using a ten-step Bayesian hyperparameter optimization scheme of Weights and Biases.